OpenStudy (anonymous):

(4a-1)(a-2)=7a-5 solve the equation check your solution show work

OpenStudy (anonymous):

one more question left after this one

OpenStudy (anonymous):

a = 1/2 a = 7/2 factored: (2 a-7) (2 a-1) = 0

OpenStudy (anonymous):

remember to visit calc-tech.com or facebook.com/calctech to download our SAT app for your math class or PSAT.

OpenStudy (anonymous):

k